Required information

[The following information applies to the questions displayed below.]

Oak Mart, a producer of solid oak tables, reports the following data from its second year of business.

  

Sales price per unit $ 320 per unit
Units produced this year 115,000 units
Units sold this year 118,000 units
Units in beginning-year inventory 3,000 units
Beginning inventory costs
Variable (3,000 units × $135) $ 405,000
Fixed (3,000 units × $80) 240,000
Total $ 645,000
Manufacturing costs this year
Direct materials $ 40 per unit
Direct labor $ 62 per unit
Overhead costs this year
Variable overhead $ 3,220,000
Fixed overhead $ 7,400,000
Selling and administrative costs this year
Variable $ 1,416,000
Fixed 4,600,000

2. Prepare the current-year income statement for the company using absorption costing.
Answer #1
2
OAK MART COMPANY
Absorption Costing Income Statement
Sales 37760000 =118000*320
Less: Cost of goods sold
Beginning inventory 645000
Manufacturing costs this year:
Direct materials 4600000 =115000*40
Direct labor 7130000 =115000*62
Variable overhead costs 3220000
Fixed overhead costs 7400000
Less: Ending inventory 0
Cost of goods sold 22995000
Gross margin 14765000
Selling general and administrative expenses
Fixed selling and administrative costs 4600000
Variable selling and administrative expenses 1416000
Total fixed expenses 6016000
Net income (loss) 8749000
Net income under variable costing is higher than net income under absorption costing by: 240,000
Number of units added to (subtracted from) inventory 3,000
Fixed overhead cost per unit 80
Fixed costs added to (subtracted from) inventory 240000